# Diffusion through complex confining environments: fluctuating triply periodic minimal surfaces

### v1

@@ V Results
Figure 3: Averaged mean squared displacement (MSD) of the diffusing spherical object over the time $t$. The three different lines are for different particle radii $l$, while $M_{\mathbf{R}}=1$ is constant. Shaded areas indicate standard deviations of the means over 64 independent realizations. Two appropriately diffusive regimes of the MSD (black line) are separated by a region of reduced slope, which is caused by trapping in the cavities of the membrane structure. With increasing particle radius, this region becomes more pronounced, for $l=8$ it is rather plateau-like.
@@
Figure 5: Averaged mean squared displacement (MSD) of the particle over the normalized time $tM_{\mathbf{R}}$. Shaded areas indicate standard deviations of the means over 64 independent realizations. The three different lines are for different mobilities $M_{\mathbf{R}}$, while $l=7$ is constant.
@@
The time axis in this figure displays the time normalized to the timescale of free diffusion, $tM_{\mathbf{R}}$, so that we can compare measurements at different particle speeds. We note that the MSD on small timescales therefore collapses on one common curve for all particle radii, as it is equal to that of a free particle. However, when the dynamics of the particle is much faster ($M_{\mathbf{R}}=10$) or much slower ($M_{\mathbf{R}}=0.1$) the long-term MSD is lower than when the two timescales coincide ($M_{\mathbf{R}}=1$). We suspect that this effect is related to the rate of jumping to the next cavity. As an underlying reason, deformation of the membrane induced by the particle makes slipping through a pore more likely. For $M_{\mathbf{R}}=10$, the membrane becomes effectively static from the point of view of the particle and does not react to its presence. In the other extreme of $M_{\mathbf{R}}=0.1$, deformations of the membrane caused by the particle relax to quickly compared to the speed of the particle.
@@
Additionally there seems to be a dynamic effect that lies in the coupling of the object and the membrane, which can speed up diffusion It goes beyond the fact that pores can open due to the fluctuations. To investigate this aspect in more detail, we introduce two more variations of the dynamics for comparison. In the first scenario the membrane is frozen in its deformed state and does not fluctuate at all. In the second variation, the membrane is not affected by the presence of the particle, while the particle is still confined by the membrane. Consequently, the system is no longer in equilibrium, as it no longer follows the stationary probability distribution Eq. (13). Nevertheless, in our theoretical approach we can simply introduce these modifications and compare the results to conclude on the underlying reasons. In Fig. 6 we compare the MSD obtained as a consequence of the two modifications to the unmodified dynamics. In a frozen environment, we find that the particle moves slower than in a fluctuating environment. This is in line with our previous results in Fig. 5, as it essentially essentially corresponds to the limit $M_{\mathbf{R}}/M_{\phi}\rightarrow\infty$. Interestingly, if one deactivates the action of the particle on the membrane, we see a comparable decrease in long-time diffusivity as well. Apparently, the effect of the particle on the membrane enhances membrane deformations, opening potential paths for the particle. It is therefore an essential part of the dynamics.
@@
Figure 6: Averaged MSD of a diffusing particle with radius $l=7$. Shaded areas indicate the standard deviation of the mean over 64 independent realizations. The full dynamics are compared to a particle diffusing through a frozen membrane, and to a membrane that fluctuates but is not deformed by the particle. In the latter two cases, the long-time MSD is significantly decreased when compared to the full dynamics.
@@
Figure 7: Long-time diffusion coefficient $D$, normalized by $D_{0}=M_{\mathbf{R}}k_{B}T$. The error bars show estimated standard deviations of the means as obtained from sampling 64 realizations. The full dynamics are compared to the scenario where we suppress the action of the particle on the membrane. For the intermediate considered value of the particle radius, the two situations lead to notably different diffusion coefficients.
We extend this comparison to different radii of the particle. In Fig. 7 we show the long time diffusion coefficient, defined by $D=\lim_{t\to\infty}\langle\mathbf{r}^{2}\rangle/6t$. It is obtained from a least squares fit to the MSD in the interval $t=10000$ to $t=50000$. Enhanced diffusion through interactions between particle and membrane is most pronounced for our intermediate radius. There is no significant difference between the dynamics with and without this coupling at $l=6$. A reason for this may be that the confinement is not strong enough in this case, as many of the pores are frequently opened much wider than the particle size. For large particles ($l=8$) the difference between the different dynamics in our observation is likewise not as pronounced as for $l=7$. One reason may be found with the comparatively high uncertainty in measurement, as only very few passings through the pores occur for large particles.
@@ VI Conclusions
In the mean squared displacement of the diffusing object, we identified two linear diffusive regimes that are separated by a plateau-like subdiffusive intermediate region that corresponds to the trapping of the particle inside a cavity of the periodic membrane structure. This region becomes broader when the size of the object is increased. The fluctuations of the membrane, particularly of the diameter of the pores, allow for the escape of a particle that would be trapped inside a given pore if the membrane surface were unperturbed. This effect should be directly observable, for instance, in protein diffusion through biological environments. Our theoretical approach allows us to separate the effect of the fluctuations of the membrane and pore size from the effect of the interaction between the diffusing object and the membrane. When by hand we suppressed deformation of the membrane by the diffusing object, the long-time diffusion coefficient decreased significantly. We concluded that the interaction between the diffusing object and the soft membrane promotes opening of the pores.
In the future, the dynamics of the phase field and the object should be coupled to the Navier-Stokes equation of hydrodynamic fluid flow to further include the role of the liquid environment [campeloDynamicModelStationary2006]. Long-range hydrodynamic interactions of the membrane with itself and with the diffusing object could be included in this way. Such couplings may play an important role in transport through biological systems as well.

### v2

@@ V Results
Figure 3: Averaged mean squared displacement (MSD) of the diffusing spherical object over the time $t$. The three different lines are for different particle radii $l$, while $M_{\mathbf{R}}=1$ is constant. Shaded areas indicate standard deviations of the means over 192 independent realizations. Two appropriately diffusive regimes of the MSD (black line) are separated by a region of reduced slope, which is caused by trapping in the cavities of the membrane structure. With increasing particle radius, this region becomes more pronounced, for $l=8$ it is rather plateau-like.
@@
Figure 5: Averaged mean squared displacement (MSD) of the particle over the normalized time $tM_{\mathbf{R}}$. Shaded areas indicate standard deviations of the means over 192 independent realizations. The three different lines are for different mobilities $M_{\mathbf{R}}$, while $l=7$ is constant.
@@
The time axis in this figure displays the time normalized to the timescale of free diffusion, $tM_{\mathbf{R}}$, so that we can compare measurements at different particle speeds. We note that the MSD on small timescales therefore collapses on one common curve for all particle radii, as it is equal to that of a free particle. However, when the dynamics of the particle is much faster ($M_{\mathbf{R}}=10$) than the dynamics of the membrane, the long-term MSD is lower than when the two timescales coincide ($M_{\mathbf{R}}=1$) or when the particle is slower ($M_{\mathbf{R}}=0.1$). We suspect that this effect is related to the rate of jumping to the next cavity. As an underlying reason, widening of pores due to thermal fluctuations makes slipping through a pore more likely. For $M_{\mathbf{R}}=10$, the membrane becomes effectively static from the point of view of the particle, so that no new pathways can open. In the other extreme of $M_{\mathbf{R}}=0.1$, deformations of the membrane relax quickly compared to the speed of the particle.
Finally, we wish to consider the role of another potential effect that may facilitate diffusion of objects through pores that, in equilibrium, are of insufficient diameter. The object itself, when pushing through its motion into the pore, may deform the membrane and widen the pore to enable passage. This process goes beyond the fact that pores can open due to the fluctuations.
@@
To investigate the relevance of these two different contributions in more detail, we introduce two variations of the dynamics for comparison. In the first scenario the membrane is frozen in its deformed state and does not fluctuate at all. The motion of the particle is confined by this frozen membrane. In the second variation, the membrane is not affected by the presence of the particle at all. That is, there is no deformational effect of the particle on the membrane. However, the particle is still confined by the membrane. Consequently, the system is no longer in equilibrium in both cases, as it no longer follows the stationary probability distribution Eq. (13). Nevertheless, in our theoretical approach we can simply introduce these modifications and compare the results to conclude on the importance of the underlying processes.
@@
Figure 6: Averaged MSD of a diffusing particle with radius $l=8$. Shaded areas indicate the standard deviation of the mean over 192 independent realizations. The full dynamics are compared to a particle diffusing through a frozen membrane, and to a membrane that fluctuates but is not deformed by the particle. In the case of a frozen membrane, the long-time MSD is significantly decreased when compared to the full dynamics.
In Fig. 6 we compare the MSD obtained from the two modifications to the unmodified equilibrium dynamics. All other parameters are kept the same, including the random seeds for the fluctuations in our simulations. In a frozen environment, we find that the particle moves slower than in a fluctuating environment. This is in line with our previous results in Fig. 5, as it essentially corresponds to the limit $M_{\mathbf{R}}/M_{\phi}\rightarrow\infty$. Interestingly, when we deactivate the action of the particle on the membrane, we see no decrease in long-time diffusivity within our standard deviations. This result confirms that the speed-up is significantly caused by the membrane fluctuations, and not notably by the force exerted by the particle on the membrane, which may contribute to widen the pores.
@@
Figure 7: Long-time diffusion coefficient $D$, normalized by $D_{0}=M_{\mathbf{R}}k_{B}T$. Error bars show estimated standard deviations from the means as obtained from sampling 192 realizations. The full dynamics are compared to the scenario of suppressed deformational action of the particle on the membrane, with no notable deviation within the standard deviations. Moreover, we compare to a scenario of a frozen membrane. The latter situation leads to notably smaller diffusion coefficients, especially for large diffusing objects.
We extend this comparison to different radii of the particle. In Fig. 7, we show the long-time diffusion coefficient, defined by $D=\lim_{t\to\infty}\langle\mathbf{r}^{2}\rangle/6t$. It is obtained from a least-squares fit to the MSD in the interval $t=20000$ to $t=100000$. With increasing size of the diffusing object, the decrease of the diffusion coefficient for a frozen membrane becomes larger. A conceivable reason refers to the confinement that is stronger for particles with larger radius, which increases the importance of fluctuation events to open the pores for passage.
@@ VI Conclusions
In the mean squared displacement of the diffusing object, we identified two linear diffusive regimes that are separated by a plateau-like subdiffusive intermediate region that corresponds to the trapping of the particle inside a cavity of the periodic membrane structure. This region becomes broader when the size of the object is increased. The fluctuations of the membrane, particularly of the diameter of the pores, allow for the escape of a particle that would be trapped inside a given pore if the membrane surface were unperturbed. This effect should be directly observable, for instance, in protein diffusion through biological environments. Our theoretical approach allows us to separate the effect of the fluctuations of the membrane and pore size from the effect that the diffusing object has on the membrane. When by hand we suppressed deformation of the membrane by the diffusing object, the long-time diffusion coefficient remained virtually unchanged. We thus conclude that the dominant effect is the opening of the pores caused by fluctuations of the membrane, and not deformations of the soft membrane induced by the diffusing object.
In the future, the dynamics of the phase field and the object should be coupled to the Navier-Stokes equation of hydrodynamic fluid flow to further include the role of the liquid environment [24]. Long-range hydrodynamic interactions of the membrane with itself and with the diffusing object could be included in this way. Such couplings may play an important role in transport through biological systems as well.